OpenStudy (gabylovesyou):

Consider the equation 7x + 3y = 42. Part 1: On your own paper, graph this equation using the slope-intercept method. In the space provided, explain, in words, each step of the procedure you used. Make sure to use complete sentences and correct grammar. (3 points) Part 2: On your own paper, graph this equation using the intercepts method. In the space provided, explain, in words, each step of the procedure you used. Make sure to use complete sentences and correct grammar. (3 points)

OpenStudy (gabylovesyou):

@chenna PLease LAST QUESTION!!

OpenStudy (gabylovesyou):

7x + 3y = 42. Part 1 and 2: On your own paper, graph this equation using the slope-intercept method. In the space provided, explain, in words, each step of the procedure you used. Make sure to use complete sentences and correct grammar. To get the equation of the line for the graph, I subtracted 7x on both sides to see what y equaled. I said that 3y equaled 42 minus 7x. I then divided by 3 on each side to get y by itself. I got y=14-7/3x for the final equation of my line. To graph this line, since 14 was the y intercept, I marked a point 14 spaces up on the y axis. From this point, since the slope is a negative slope and slope is equivalent to rise divided by run on a graph, I went up 7 spaces then 3 spaces to the left, then marked a point there, then I finally connected the two points on the graph I made to draw my line. For using the slope method, the two points I got were (3,7) and (0,14). I used the formula y2-y1/x2-x1 and solved for the slope. (14-7)/(0-3) (7)/(-3) -7/3 is still the slope I got for this line. Then with the points (3,7) and the slope, I plugged those points in the point slope form equation to check my work. y-y1=m(x-x1) y-7=-7/3(x-3) y-7=-7/3x+7 y=-7/3x+14 is still the equation of the line I got for the graph.

OpenStudy (gabylovesyou):

@chenna is that ok?? ^^

OpenStudy (anonymous):

no. you need to draw the graph and explain how you drew it .